Which fraction is equivalent to
1 1/2 - 7/8
A.5/12
B.5/96
C.5/48
D.5/24

Answer by math_helper(2461)

Which fraction is equivalent to
1 1/2 - 7/8
A.5/12
B.5/96
C.5/48
D.5/24
========================================
None of the above.
1 1/2 - 7/8 = 3/2 - 7/8 = 12/8 - 7/8 = 5/8, which is not equivalent to any of the choices.